OpenStudy (anonymous):

Nine less than 4 times a number is the same as twice the number increased by 11. What is the unknown number?

OpenStudy (anonymous):

10 12 18 23

OpenStudy (anonymous):

4x - 9 = 2x+11 4x - 9 = 2x + 11 2x = 20 x = 10